OpenStudy (anonymous):

What is the volume of the oblique square pyramid with base edges 25ft and height 24 ft?

OpenStudy (anonymous):

1/3*25*25*24\[5000ft ^{3}\]

OpenStudy (anonymous):

V = (1/3)*B*h B = area of base h = height B(square) = 25ft * 25ft = 625ft^2 V = (1/3)*625ft^2*24ft V = 5000ft^3